Find the value of c so that the line passing through (1,2) and (3,c) is parallel to y= -3x+5/2.
Answer by mananth(16946)

(1,2) and (3,c) is parallel to


y= -3x+5/2.
slope m = -3
slope between two points
-3= (c-2)/(3-1)
-3*2=c-2
-6=c-2
c=-4